A 1 by 1 meter mirror reflects light with a power of 80 W and passes through a polarizing film whose axis is 0o from vertical. The liquid crystal and the 2nd polarizing film combined axis is 45o from the horizontal. What is the power of the light as it emerges from the 2nd film? Assume voltage is held constant.
Answer #1

Assuming the incident light is unpolarized, the power after the first polarizing film will be:

P = Po/2 = 40 W

now, use Malus' law

P' = Pcos^2\theta

therefore, P' = 40cos^245\degree = 20 W

this is the power of the light as it emerges from the 2nd film.